OpenStudy (anonymous):

A group decides to go whitewater rafting on a river. The map they have of the river has the scale 3 inches: 7 miles. On the map, the distance between the starting point and the endpoint of their trip is 15 inches. What is the actual distance? A. 2.33 miles B. 49 miles C. 6.43 miles D. 35 miles its d right

OpenStudy (anonymous):

proportion time.. 3/7 = 15/x (3 in to 7 miles = 15 in to x miles) cross multiply (3)(x) = (15)(7) 3x = 105 -- divide by 3 x = 105/3 x = 35 miles D is correct :)

OpenStudy (anonymous):

thank you! @kelliegirl33

OpenStudy (anonymous):

no problem :)